Suppose four firms have market shares of 30%, 30%, 20% and 20%. What is the Herfindahl-Hirschman Index for this market?

10000

2600

100

25

Question 18

What is the Herfindahl–Hirschman Index for a market with 100 firms, each with 1% market share?

10000

1000

100

1
Answer #1

Answer : 1) The answer is option B.

Herfindahl-Hirschman Index (HHI) = The sum of all available firms' market share's square

=> HHI = (30)^2 + (30)^2 + (20)^2 + (20)^2

=> HHI = 900 + 900 + 400 + 400

=> HHI = 2600

Therefore, option B is correct.

18) The answer is option C.

HHI = Square of market share * number of firms [ As here each firm's market share is 1%]

=> HHI = (1)^2 * 100

=> HHI = 100

Therefore, option C is correct.